OpenStudy (anonymous):

a merchant has coffee worth $60 a pound and she wants to mix with 70 pounds of coffee worth $90 a pound to get a mixture that is worth $70a pound. How many pounds of $60 coffee should be used. A.105lb B. 210lb C. 70lb D. 140lb

OpenStudy (anonymous):

you have \(90\times 70=6300\) worth of coffee at present. if you add \(x\) amount of $60 coffee you will have \(6300+60x\) worth of coffee and a total of \(70+x\) pounds of coffee that you want to be worth $70 a pound, so set \[6300+60x=70(70+x)\] and solve for \(x\)

OpenStudy (anonymous):

ok working it now

OpenStudy (anonymous):

140